# Dual metrics on the boundary of strictly polyhedral hyperbolic 3-manifolds

### v4

@@ 1.1 Metric geometry of convex polyhedra
One can see that the conditions on $d$ in Theorem 1.3 are also necessary (the only non-trivial condition is largeness, and its necessity is proven in [56]). Theorem 1.3 can be seen as a generalization of the Andreev Theorem [3] on compact convex polyhedra in $\mathbb{H}^{3}$ with non-obtuse dihedral angles, another generalization of which was famously used by Thurston in his Hyperbolization Theorem for Haken 3-manifolds, see [66, 46]. Theorem 1.3 was also used by Rivin in [54] to describe the dihedral angles of ideal polyhedra in $\mathbb{H}^{3}$, which in turn implied a resolution of the Steiner Problem. The latter asked how to distinguish the combinatorial types of convex Euclidean 3-polyhedra that admit a realization with all vertices on the standard 2-sphere. In [58] Schlenker proved a version of Theorem 1.3 for smooth strictly convex bodies in $\mathbb{H}^{3}$. We note that if $C\subset\mathbb{H}^{3}$ is a smooth strictly convex body, then the pull-back of the dual metric by the Gauss map is exactly the third fundamental form of $\partial C$.
@@ 1.2 Hyperbolic 3-manifolds with convex boundary
In the seventies Thurston revolutionized the field of 3-dimensional topology by showing the importance of geometric structures on 3-manifolds, posing his famous Geometrization Conjecture (whose proof was completed by Perelman) and demonstrating a special role of hyperbolic geometry in dimension 3. In this paper we will be interested in studying compact hyperbolic 3-manifolds with convex polyhedral boundary.
@@
The proof is decomposed into three pieces: establishing differentiability properties of $\mathcal{I}_{V}$ and its infinitesimal rigidity, showing its properness, and studying connectivity properties of the spaces $\mathcal{P}(M,V)$ and $\mathcal{D}_{M}(\partial M,V)$. The last item was basically done before us: it is a generalization of the connectivity results from [56], as it was mentioned by Schlenker in [61]. Thus, we mostly focus on the infinitesimal rigidity and the properness. Note also that the connectivity results on the spaces above are obtained only in some weak sense, and a special topological argument is required to finish the proof.
@@
Up to now we discussed in parallel questions for polyhedral and for smooth boundaries. It is very natural to ask if there is a common generalization unifying these two settings. This was famously initiated by Alexandrov in his work on general metrics on surfaces with curvature bounded from below [1]. He particularly proved a realization theorem, obtaining a complete intrinsic classification of the induced path metrics on the boundaries of convex bodies in $\mathbb{E}^{3}$, $\mathbb{H}^{3}$ or $\mathbb{S}^{3}$. The rigidity question for general convex bodies was settled by Pogorelov [49] and is known as notoriously difficult. In recent years there was a stream of work showing general realization results on hyperbolic 3-manifolds, e.g., [63, 24]. The only known general rigidity result for manifolds topologically more complicated than the 3-ball (for Fuchsian manifolds of the first kind) was done by the author in [50]. The corresponding results for the dual metrics are still not known. An important particular case, again, is given by the convex cores of convex cocompact hyperbolic 3-manifolds. Their dual metric is somewhat degenerate, but can be replaced with another object called the pleating lamination measuring how their boundary is bent. Thurston conjectured that convex cores are uniquely determined by their pleating lamination as well as by the induced boundary metric. The realization part of the former question was shown in [6] and [35] by Bonahon–Otal and Lecuire, while for the latter it is, e.g., a particular corollary of the work [34] by Labourie, where another proof of the realization part of Theorem 1.6 was given. The rigidity part remains to be an open problem. This is reminiscent to the famous Ahlfors–Bers Theorem (see Theorem 2.11), which could be stated as that convex cocompact hyperbolic metrics on 3-manifolds are determined up to isotopy by the conformal structure at infinity.
Acknowledgments. We thank François Fillastre, Ivan Izmestiev, Jean-Marc Schlenker and the anonymous referees for helpful comments, in particular, leading to a simplification of the proof of Lemma 2.17. We are also grateful to our mentor, Michael Eichmair, for his support.
@@ 2.1 Hyperbolic and de Sitter 3-spaces and the duality
endowed with the induced metric. It is a Lorentzian 3-manifold of constant curvature 1 homeomorphic to $S^{2}\times\mathbb{R}$. Let us say that a subset of $\mathbb{H}^{3}$ or $d\mathbb{S}^{3}$ is properly convex when it is proper, has non-empty interior and is the intersection of closed half-spaces (the latter is equivalent to be closed and convex in $\mathbb{H}^{3}$, but is more subtle in $d\mathbb{S}^{3}$).
@@ 2.2 Cone-metrics
We say that a sequence of metrics $\{d_{i}\}$ on a surface $S$ converges to a metric $d$ in the Lipschitz sense if the identity maps $(S,d_{i})\rightarrow(S,d)$ have distortions $\varepsilon_{i}\rightarrow 0$. We remark that some authors allow the action by homeomorphisms in the definitions of the Lipschitz convergence, e.g., in [56, Section 6]. In a more general setting the Lipschitz convergence may be defined between non-homeomorphic metric spaces, see [9, Section 7.2]. We denote by $\mathfrak{D}(S,V)$ the set of spherical cone-metrics on $(S,V)$ that are $\mathcal{T}$-triangulable for at least one triangulation. We endow it with the topology of Lipschitz convergence.
@@
Pick $d\in\mathfrak{D}(S,\mathcal{T})$. There is a canonical section through $d$ of the forgetful map $\mathfrak{D}(S,\mathcal{T})\rightarrow\mathcal{D}^{\sharp}(S,\mathcal{T})$. Indeed, for any two convex spherical triangles $T_{1}$ and $T_{2}$ with marked vertices there exists a canonical “affine” homeomorphism between them preserving the marking. Indeed, develop $T_{1}$ and $T_{2}$ on the standard sphere $\mathbb{S}^{2}$ embedded in Euclidean 3-space $\mathbb{R}^{3}$. Project $T_{1}$ radially to the Euclidean triangle subtending it, send it by the unique linear map to the affine triangle subtending $T_{2}$ and project it radially to $T_{2}$. This defines the desired map. For every class $d\in\mathfrak{D}(S,\mathcal{T})$ such affine maps produce the canonical section of $\mathfrak{D}(S,\mathcal{T})\rightarrow\mathcal{D}^{\sharp}(S,\mathcal{T})$ through $d$. Acting by $H^{\sharp}_{0}(S,V)$ on the metrics of this section, one can see that $\mathfrak{D}(S,\mathcal{T})\cong\mathcal{D}^{\sharp}(S,\mathcal{T})\times H^{\sharp}_{0}(S,V)$. This produces the structure of a principal bundle on $\mathfrak{D}(S,V)\rightarrow\mathcal{D}^{\sharp}(S,V)$ with the structure group $H^{\sharp}_{0}(S,V)$. Since the latter is known to be contractible [28], this bundle is trivial.
@@
We denote by $\mathbb{S}^{1}_{l}$ the metric circle of length $l$. A closed global geodesic in a metric space $(S,d)$ is the image of an isometric map $\mathbb{S}^{1}_{l}\rightarrow(S,d)$ (here it is crucial that it preserves distances in $\mathbb{S}^{1}_{l}$ globally). A closed $\varepsilon$-quasi-geodesic is the image of $S^{1}_{l}$ under a map $S^{1}_{l}\rightarrow(S,d)$ with distortion at most $\varepsilon$. The following consequence of the Arzelá–Ascoli Theorem was shown in [56, Theorem 6.1]:
@@ 2.3 Convex cocompact hyperbolic manifolds
Let $g$ be a hyperbolic metric on $M$ such that $\partial M$ is locally convex. Recall that we denote by $G$ the group of orientation-preserving isometries of $\mathbb{H}^{3}$. It is naturally isomorphic to $PSL(2,\mathbb{C})$.
@@
Abusing the notation, we denote the extension of $\overline{g}$ to the universal cover $\widetilde{N}$ still by $\overline{g}$, and the hyperbolic 3-manifold $(\widetilde{N},\overline{g})$ by $\widetilde{N}(\overline{g})$. We will frequently identify $\widetilde{N}(\overline{g})$ with $\mathbb{H}^{3}$ by an arbitrary developing map (recall that a developing map is defined up to a global isometry of $\mathbb{H}^{3}$). We denote by $\widetilde{M}(g)$ the respective embedding of the universal cover of $M(g)$ as a convex set. Let $\rho_{\overline{g}}:\pi_{1}(M)\rightarrow G$ be the corresponding holonomy representation. We denote by $\Lambda(\overline{g})=\Lambda(\rho_{\overline{g}})$ its limit set and by $\widetilde{C}(\overline{g})=\widetilde{C}(\rho_{\overline{g}})$ we denote the convex hull of $\Lambda(\overline{g})$. If $\Lambda(\overline{g})$ is contained in a circle, we say that $\overline{g}$ and $\rho_{\overline{g}}$ are Fuchsian. They are Fuchsian of the first kind if $\Lambda(\overline{g})$ is a circle, and of the second kind otherwise. The set $\widetilde{C}(\overline{g})$ is $\rho_{\overline{g}}$-invariant and projects to the convex core $C(\overline{g})\subset N$. The latter is totally convex, and is homeomorphic to $M$, except in the Fuchsian case. We will refer to the connected components of $N(\overline{g})\backslash C(\overline{g})$ and of $\widetilde{N}(\overline{g})\backslash\widetilde{C}(\overline{g})$ as to the geometric ends of $N(\overline{g})$ and $\widetilde{N}(\overline{g})$. Since the $\rho_{\overline{g}}$-orbit of any point of $\mathbb{H}^{3}$ accumulates at $\Lambda(\overline{g})$, $C(\overline{g})$ is contained in any totally convex subset of $N(\overline{g})$. It is also well-known that
@@
Denote by $\mathcal{CH}(N)$ the space of convex cocompact hyperbolic metrics on $N$ up to isotopy. The rest of this section is devoted to the discussion of a topology on this space. While it seems to be folklore, we did not find in the literature a reference suitable to our needs. Thereby, we devote some time to extract it from the existing literature. We will frequently abuse the notation and write $\overline{g}\in\mathcal{CH}(N)$ meaning that $\overline{g}$ is a concrete convex cocompact hyperbolic metric on $N$.
@@
We denote by $\mathcal{R}(\pi_{1}(M),G)$ the space of representations of $\pi_{1}(M)$ in $G$. The group $G$ acts on $\mathcal{R}(\pi_{1}(M),G)$ by conjugation, and we denote the resulting quotient by $\mathcal{X}(\pi_{1}(M),G)$. Since $G\cong PSL(2,\mathbb{C})$ is an algebraic group over $\mathbb{C}$, the space $\mathcal{R}(\pi_{1}(M),G)$ can be viewed as a complex algebraic variety, and we consider it with the induced topology. We endow the space $\mathcal{X}(\pi_{1}(M),G)$ with the quotient topology. Note that $\mathcal{X}(\pi_{1}(M),G)$ is not even a Hausdorff space, and should not be confused with the character variety, which is a more accurate quotient that can be constructed by means of algebraic geometry. However, we are interested only in the quotient of an invariant open smooth subset of $\mathcal{R}(\pi_{1}(M),G)$, corresponding to the holonomies of convex cocompact metrics, over which the action behaves well. We will be anyway calling the elements of $\mathcal{X}(\pi_{1}(M),G)$ characters.
@@
Let $\overline{g}\in\mathcal{CH}(N)$ and $\rho_{\overline{g}}$ be its holonomy. It follows from Marden’s Isomorphism Theorem [36, Section 8] that a convex cocompact representation $\rho\in\mathcal{R}(\pi_{1}(M),G)$ is a holonomy representation of a convex cocompact metric on $N$ if and only if it is related to $\rho_{\overline{g}}$ by a quasi-conformal deformation. Thus, the set of holonomy representations of metrics from $\mathcal{CH}_{h}(N)$ can be identified with the set $\mathcal{QC}(\rho_{\overline{g}})$ of quasi-conformal deformations of $\rho_{\overline{g}}$. In particular, it is connected. Marden’s Stability Theorem [36, Section 10] implies also that it is an open subset of the smooth part of $\mathcal{R}(\pi_{1}(M),G)$. The set $\mathcal{QC}(\rho_{\overline{g}})$ stays invariant under the action of $G$ by conjugation, and $G$ acts freely and smoothly on it. Moreover, [36, Theorem 10.8] implies that this action is proper, so the quotient is a complex manifold. Other approaches to the properness can be seen in [32, Proposition 1.1] or in [70, Lemma 6.24]. We identify $\mathcal{CH}_{h}(N)$ with this quotient. We refer to the expositions in [38, Chapter 5] and [11, Section 7].
@@
The work of Waldhausen [68] establishes that when the boundary of $M$ is incompressible, the group $\textrm{MCG}_{h}(G)$ is trivial. However, this is no longer the case when the boundary is compressible. Let $\gamma\subset\partial M$ be an oriented simple closed curve that is contractible in $M$, and $D\subset M$ be an embedded 2-disk contracting $\gamma$, which exists thanks to the Dehn Lemma. Do a Dehn twist around $\gamma$. Let $D\times[-1,1]\subset M$ be a small neighborhood of $D$ such that $D\times\{0\}=D$ and $\partial D\times[-1,1]$ is an annulus around $\gamma$ supporting the Dehn twist. The Dehn twist naturally extends to $D\times[-1,1]$. Extending it by identity outside of $D\times[-1,1]$, we obtain a self-homeomorphism of $M$ called a twist of $M$. One can see that it is homotopic to the identity, but not isotopic. McCullough and Miller showed in [41] that the kernel of $\textrm{MCG}(M)\rightarrow{\rm Out}(\pi_{1}(M))$ is generated by twists. Thereby, $\textrm{MCG}_{h}(M)$ is equal to the kernel and is generated by twists. Let $S_{j}$ be a connected component of $\partial M$. We denote by $\textrm{MCG}_{h,j}(M)$ the subgroup of $\textrm{MCG}_{h}(M)$ generated by twists along curves belonging to $S_{j}$. As compressing disks of disjoint simple closed curves in $\partial M$ can be made disjoint, $\textrm{MCG}_{h}(M)$ is the direct product of all $\textrm{MCG}_{h,j}(M)$.
We denote by $\mathcal{TS}(\partial M)$ the Teichmüller space of $\partial M$. It is the direct product of the Teichmüller spaces of all connected components of $\partial M$. The group $\textrm{MCG}_{h}(M)$ acts on $\mathcal{TS}(\partial M)$ and the action splits on actions of $\textrm{MCG}_{h,j}(M)$ on the respective components of $\mathcal{TS}(\partial M)$. As the action of the mapping class group of $\partial M$ on $\mathcal{TS}(\partial M)$ is propely discontinuous (see, e.g., [18, Theorem 12.2]), so is the action of $\textrm{MCG}_{h}(M)$. In [37, Section 3.1] Marden shows that this action is free. We denote by $\mathcal{TS}_{h}(\partial M,M)$ the quotient of $\mathcal{TS}(\partial M)$ over this action.
In [37, Section 3.1] Marden states a version of the Ahlfors–Bers Theorem (based also on the works of Kra and Maskit), which says that if $\overline{g}$ is a convex cocompact hyperbolic metric on $N$, then the action of quasi-conformal deformations of $\rho_{\overline{g}}$ on $\partial_{\infty}\mathbb{H}^{3}$ defines a biholomorphic map from the quotient of $\mathcal{QC}(\rho_{\overline{g}})$ by conjugation of $G$ to $\mathcal{TS}_{h}(\partial M,M)$. We refer to [37] and references therein. Additional expositions can be found in [38, Section 5.1] or [11, Section 7]. Due to the discussion above, we can state this result as

### v5

@@ 1.1 Metric geometry of convex polyhedra
One can see that the conditions on $d$ in Theorem 1.3 are also necessary (the only non-trivial condition is largeness, and its necessity is proven in [58]). Theorem 1.3 can be seen as a generalization of the Andreev Theorem [3] on compact convex polyhedra in $\mathbb{H}^{3}$ with non-obtuse dihedral angles, another generalization of which was famously used by Thurston in his Hyperbolization Theorem for Haken 3-manifolds, see [69, 48]. For a corrected proof of the Andreev Theorem and more references we refer to [59]. Theorem 1.3 was also used by Rivin in [56] to describe the dihedral angles of ideal polyhedra in $\mathbb{H}^{3}$, which in turn implied a resolution of the Steiner Problem. The latter asked how to distinguish the combinatorial types of convex Euclidean 3-polyhedra that admit a realization with all vertices on the standard 2-sphere. A more general Steiner Problem, with the 2-sphere replaced by the other types of quadrics in $\mathbb{RP}^{3}$, was resolved in [17]. In [61] Schlenker proved a version of Theorem 1.3 for smooth strictly convex bodies in $\mathbb{H}^{3}$. We note that if $C\subset\mathbb{H}^{3}$ is a smooth strictly convex body, then the pull-back of the dual metric by the Gauss map is exactly the third fundamental form of $\partial C$.
@@ 1.2 Hyperbolic 3-manifolds with convex boundary
In the seventies Thurston revolutionized the field of 3-dimensional topology by showing the importance of homogeneous geometric structures on 3-manifolds, posing his famous Geometrization Conjecture (whose proof was completed by Perelman) and demonstrating a special role of hyperbolic geometry in dimension 3. In this paper we will be interested in studying compact hyperbolic 3-manifolds with convex polyhedral boundary.
@@
The proof is decomposed into three pieces: establishing differentiability properties of $\mathcal{I}_{V}$ and its infinitesimal rigidity, showing its properness, and studying connectivity properties of the spaces $\mathcal{P}(M,V)$ and $\mathcal{D}_{M}(\partial M,V)$. The last item was basically done before us: it is a generalization of the connectivity results from [58], as was mentioned by Schlenker in [64]. Thus, we mostly focus on the infinitesimal rigidity and the properness. Note also that the connectivity results on the spaces above are obtained only in some weak sense, and a special topological argument is required to finish the proof.
@@
Up to now we discussed in parallel questions for polyhedral and for smooth boundaries. It is very natural to ask if there is a common generalization unifying these two settings. This was famously initiated by Alexandrov in his work on general metrics on surfaces with curvature bounded from below [1]. He particularly proved a realization theorem, obtaining a complete intrinsic classification of the induced path metrics on the boundaries of convex bodies in $\mathbb{E}^{3}$, $\mathbb{H}^{3}$ or $\mathbb{S}^{3}$. The rigidity question for general convex bodies was settled by Pogorelov [51] and is known as notoriously difficult. In recent years there was a stream of work showing general realization results on hyperbolic 3-manifolds, e.g., [66, 26]. The only known general rigidity result for manifolds topologically more complicated than the 3-ball (for Fuchsian manifolds of the first kind) was done by the author in [52]. The corresponding results for the dual metrics are still not known. An important particular case, again, is given by the convex cores of convex cocompact hyperbolic 3-manifolds. Their dual metric is somewhat degenerate, but can be replaced with another object called the pleating lamination measuring how their boundary is bent. Thurston conjectured that convex cores are uniquely determined by their pleating lamination as well as by the induced boundary metric. The realization part of the former question was shown in [6] and [37] by Bonahon–Otal and Lecuire, while for the latter it is, e.g., a particular corollary of the work [36] by Labourie, where another proof of the realization part of Theorem 1.6 was given. The rigidity part for pleating laminations was recently resolved by Dular–Schlenker in [18], while for the induced metric it remains an open problem, with some progress obtained in [54]. This is reminiscent to the famous Ahlfors–Bers Theorem (see Theorem 2.12), which could be stated as that convex cocompact hyperbolic metrics on 3-manifolds are determined up to isotopy by the conformal structure at infinity.
Acknowledgments. We thank François Fillastre, Ivan Izmestiev, Jean-Marc Schlenker and the anonymous referees for helpful comments. We are also grateful to our mentor, Michael Eichmair, for his support.
@@ 2.1 Hyperbolic and de Sitter 3-spaces and the duality
endowed with the induced metric. It is a Lorentzian 3-manifold of constant curvature 1 homeomorphic to $S^{2}\times\mathbb{R}$. Let us say that a subset of $\mathbb{H}^{3}$ or $d\mathbb{S}^{3}$ is properly convex when it is proper, has non-empty interior and is the intersection of closed half-spaces (the latter is equivalent to be closed and convex in $\mathbb{H}^{3}$, but the convexity is more subtle in $d\mathbb{S}^{3}$).
@@ 2.2 Cone-metrics
We say that a sequence of metrics $\{d_{i}\}$ on a surface $S$ converges to a metric $d$ in the Lipschitz sense if the identity maps $(S,d_{i})\rightarrow(S,d)$ have distortions $\varepsilon_{i}\rightarrow 0$. We remark that some authors allow the action by homeomorphisms in the definitions of the Lipschitz convergence, e.g., in [58, Section 6]. In a more general setting the Lipschitz convergence may be defined between non-homeomorphic metric spaces, see [9, Section 7.2]. We denote by $\mathfrak{D}(S,V)$ the set of spherical cone-metrics on $(S,V)$ that are $\mathcal{T}$-triangulable for at least one triangulation. We endow it with the topology of Lipschitz convergence.
@@
Pick $d\in\mathfrak{D}(S,\mathcal{T})$. There is a canonical section through $d$ of the forgetful map $\mathfrak{D}(S,\mathcal{T})\rightarrow\mathcal{D}^{\sharp}(S,\mathcal{T})$. Indeed, for any two convex spherical triangles $T_{1}$ and $T_{2}$ with marked vertices there exists a canonical “affine” homeomorphism between them preserving the marking. To see this, develop $T_{1}$ and $T_{2}$ on the standard sphere $\mathbb{S}^{2}$ embedded in Euclidean 3-space $\mathbb{R}^{3}$. Project $T_{1}$ radially to the Euclidean triangle subtending it, send it by the unique linear map to the affine triangle subtending $T_{2}$ and project it radially to $T_{2}$. This defines the desired map. For every class $d\in\mathfrak{D}(S,\mathcal{T})$ such affine maps produce the canonical section of $\mathfrak{D}(S,\mathcal{T})\rightarrow\mathcal{D}^{\sharp}(S,\mathcal{T})$ through $d$. Acting by $H^{\sharp}_{0}(S,V)$ on the metrics of this section, one can see that $\mathfrak{D}(S,\mathcal{T})\cong\mathcal{D}^{\sharp}(S,\mathcal{T})\times H^{\sharp}_{0}(S,V)$. This produces the structure of a principal bundle on $\mathfrak{D}(S,V)\rightarrow\mathcal{D}^{\sharp}(S,V)$ with the structure group $H^{\sharp}_{0}(S,V)$. Since the latter is known to be contractible [30], this bundle is trivial.
@@
We denote by $\mathbb{S}^{1}_{l}$ the metric circle of length $l$. A closed global geodesic in a metric space $(S,d)$ is the image of an isometric map $\mathbb{S}^{1}_{l}\rightarrow(S,d)$ (here it is crucial that it preserves distances in $\mathbb{S}^{1}_{l}$ globally). A closed $\varepsilon$-quasi-geodesic is the image of $S^{1}_{l}$ under a map $S^{1}_{l}\rightarrow(S,d)$ with distortion at most $\varepsilon$. The following consequence of the Arzelá–Ascoli Theorem was shown in [58, Theorem 6.1]:
@@ 2.3 Convex cocompact hyperbolic manifolds
Let $g$ be a hyperbolic metric on $M$ such that $\partial M$ is locally convex. Recall that we denote by $G$ the group of orientation-preserving isometries of $\mathbb{H}^{3}$. It is naturally isomorphic to ${\rm PSL}(2,\mathbb{C})$.
@@
Abusing the notation, we denote the extension of $\overline{g}$ to the universal cover $\widetilde{N}$ still by $\overline{g}$, and the hyperbolic 3-manifold $(\widetilde{N},\overline{g})$ by $\widetilde{N}(\overline{g})$. We will frequently identify $\widetilde{N}(\overline{g})$ with $\mathbb{H}^{3}$ by an arbitrary developing map (recall that a developing map is defined up to a global isometry of $\mathbb{H}^{3}$). We denote by $\widetilde{M}(g)$ the respective embedding of the universal cover of $M(g)$ as a convex set. Let $\rho_{\overline{g}}:\pi_{1}(M)\rightarrow G$ be the corresponding holonomy representation. We denote by $\Lambda(\overline{g})=\Lambda(\rho_{\overline{g}})$ its limit set and by $\widetilde{C}(\overline{g})=\widetilde{C}(\rho_{\overline{g}})$ we denote the convex hull of $\Lambda(\overline{g})$. If $\Lambda(\overline{g})$ is contained in a circle, we say that $\overline{g}$ and $\rho_{\overline{g}}$ are Fuchsian. They are Fuchsian of the first kind if $\Lambda(\overline{g})$ is a circle, and of the second kind otherwise. The set $\widetilde{C}(\overline{g})$ is $\rho_{\overline{g}}$-invariant and projects to the convex core $C(\overline{g})\subset N(\overline{g})$. The latter is totally convex, and is homeomorphic to $M$, except in the Fuchsian case. We will refer to the connected components of $N(\overline{g})\backslash C(\overline{g})$ and of $\widetilde{N}(\overline{g})\backslash\widetilde{C}(\overline{g})$ as to the geometric ends of $N(\overline{g})$ and $\widetilde{N}(\overline{g})$. Since the $\rho_{\overline{g}}$-orbit of any point of $\mathbb{H}^{3}$ accumulates at $\Lambda(\overline{g})$, $C(\overline{g})$ is contained in any totally convex subset of $N(\overline{g})$. It is also well-known that
@@
Denote by $\mathcal{CH}(N)$ the space of convex cocompact hyperbolic metrics on $N$, where we consider two metrics equivalent if they are related by a self-homeomorphism of $M$ isotopic to the identity on $M$, where we use $N\cong{\rm int}(M)$. The rest of this section is devoted to the discussion of a topology on this space. While it seems to be folklore, we did not find in the literature a reference suitable to our needs. Thereby, we devote some time to extract it from the existing literature. We will frequently abuse the notation and write $\overline{g}\in\mathcal{CH}(N)$ meaning that $\overline{g}$ is a concrete convex cocompact hyperbolic metric on $N$. Note that one can also consider metrics up to isotopy on $N$, which a priori can lead to a stricter equivalence. However, this is not the case, see Remark 2.10, but we will not use it.
@@
We denote by $\mathcal{R}(\pi_{1}(M),G)$ the space of representations of $\pi_{1}(M)$ in $G$. The group $G$ acts on $\mathcal{R}(\pi_{1}(M),G)$ by conjugation, and we denote the resulting quotient by $\mathcal{X}(\pi_{1}(M),G)$. Since $G\cong{\rm PSL}(2,\mathbb{C})$ is an algebraic group over $\mathbb{C}$, the space $\mathcal{R}(\pi_{1}(M),G)$ can be viewed as a complex algebraic variety, and we consider it with the induced topology. We endow the space $\mathcal{X}(\pi_{1}(M),G)$ with the quotient topology. Note that $\mathcal{X}(\pi_{1}(M),G)$ is not even a Hausdorff space, and should not be confused with the character variety, which is a more accurate quotient that can be constructed by means of algebraic geometry. However, we are interested only in the quotient of an invariant open smooth subset of $\mathcal{R}(\pi_{1}(M),G)$, corresponding to the holonomies of convex cocompact metrics, over which the action behaves well. We will be anyway calling the elements of $\mathcal{X}(\pi_{1}(M),G)$ characters.
@@
Let $\overline{g}\in\mathcal{CH}_{h}(N)$ and $\rho_{\overline{g}}$ be its holonomy. It follows from Marden’s Isomorphism Theorem [38, Section 8] that a convex cocompact representation $\rho\in\mathcal{R}(\pi_{1}(M),G)$ is a holonomy representation of a convex cocompact metric on $N$ if and only if it is related to $\rho_{\overline{g}}$ by a quasi-conformal deformation. Thus, the set of holonomy representations of metrics from $\mathcal{CH}_{h}(N)$ can be identified with the set $\mathcal{QC}(\rho_{\overline{g}})$ of quasi-conformal deformations of $\rho_{\overline{g}}$. In particular, it is connected. Marden’s Stability Theorem [38, Section 10] implies also that it is an open subset of the smooth part of $\mathcal{R}(\pi_{1}(M),G)$. The set $\mathcal{QC}(\rho_{\overline{g}})$ stays invariant under the action of $G$ by conjugation, and $G$ acts freely and smoothly on it. Moreover, [38, Theorem 10.8] implies that this action is proper, so the quotient is a complex manifold. Other approaches to the properness can be seen in [34, Proposition 1.1] or in [73, Lemma 6.24]. We identify $\mathcal{CH}_{h}(N)$ with this quotient. We refer to the expositions in [40, Chapter 5] and [11, Section 7].
@@
The work of Waldhausen [71] establishes that when the boundary of $M$ is incompressible, the group $\textrm{MCG}_{h}(M)$ is trivial. However, this is no longer the case when the boundary is compressible. Let $\gamma\subset\partial M$ be an oriented simple closed curve that is contractible in $M$, and $D\subset M$ be an embedded 2-disk contracting $\gamma$, which exists thanks to the Dehn Lemma. Do a Dehn twist around $\gamma$. Let $D\times[-1,1]\subset M$ be a small neighborhood of $D$ such that $D\times\{0\}=D$ and $\partial D\times[-1,1]$ is an annulus around $\gamma$ supporting the Dehn twist. The Dehn twist naturally extends to $D\times[-1,1]$. Extending it by identity outside of $D\times[-1,1]$, we obtain a self-homeomorphism of $M$ called a twist of $M$. One can see that it is homotopic to the identity, but not isotopic. McCullough and Miller showed in [43] that the kernel of $\textrm{MCG}(M)\rightarrow{\rm Out}(\pi_{1}(M))$ is generated by twists. Thereby, $\textrm{MCG}_{h}(M)$ is equal to the kernel and is generated by twists. Let $S_{j}$ be a connected component of $\partial M$. We denote by $\textrm{MCG}_{h,j}(M)$ the subgroup of $\textrm{MCG}_{h}(M)$ generated by twists along curves belonging to $S_{j}$. As compressing disks of disjoint simple closed curves in $\partial M$ can be made disjoint, $\textrm{MCG}_{h}(M)$ is the direct product of all $\textrm{MCG}_{h,j}(M)$.
Remark 2.10.
One can consider a similarly-defined group $\textrm{MCG}_{h}(N)$. We have a morphism $\textrm{MCG}_{h}(M)\rightarrow\textrm{MCG}_{h}(N)$. One can show that this is an isomorphism. We will not rely on it, hence we do not go into the details of it.
We denote by $\mathcal{TS}(\partial M)$ the Teichmüller space of $\partial M$. It is the direct product of the Teichmüller spaces of all connected components of $\partial M$. The group $\textrm{MCG}_{h}(M)$ acts on $\mathcal{TS}(\partial M)$ and the action splits on actions of $\textrm{MCG}_{h,j}(M)$ on the respective components of $\mathcal{TS}(\partial M)$. As the action of the mapping class group of $\partial M$ on $\mathcal{TS}(\partial M)$ is propely discontinuous (see, e.g., [20, Theorem 12.2]), so is the action of $\textrm{MCG}_{h}(M)$. In [39, Section 3.1] Marden shows that this action is free. We denote by $\mathcal{TS}_{h}(\partial M,M)$ the quotient of $\mathcal{TS}(\partial M)$ over this action.
In [39, Section 3.1] Marden states a version of the Ahlfors–Bers Theorem (based also on the works of Kra and Maskit), which says that if $\overline{g}$ is a convex cocompact hyperbolic metric on $N$, then the action of quasi-conformal deformations of $\rho_{\overline{g}}$ on $\partial_{\infty}\mathbb{H}^{3}$ defines a biholomorphic map from the quotient of $\mathcal{QC}(\rho_{\overline{g}})$ by conjugation of $G$ to $\mathcal{TS}_{h}(\partial M,M)$. We refer to [39] and references therein. Additional expositions can be found in [40, Section 5.1] or [11, Section 7]. Due to the discussion above, we can state this result as